OpenStudy (anonymous):

how can i factor out this completely: 28b^2-80b-128?

OpenStudy (anonymous):

What is the GCF of 28, 80 and 128?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Factor out 4 4(7b^2 - 20b -32) What multiplies to make 7? What multiples to make 32? <------ be careful

OpenStudy (anonymous):

1*7 1*32,2*16,4*8

OpenStudy (anonymous):

but that's where i'm stuck!!!

OpenStudy (anonymous):

look back at the last question. I am not sure you know why they added to -38. I think you might have hurried over that part. Maybe I am wrong? For 4(7b^2 - 20b -32) we have 4(7b ?)(b ?) which ones should we choose for 32?

OpenStudy (anonymous):

i don't know....

OpenStudy (anonymous):

We need to get -20 If we take 4 and 8 notice 7 * 4 = 28 1 * 8 = 8 Notice 28 - 8 = 20 ^ ^ | | | | | | | multiplies to make 32 | multiples to make 7

OpenStudy (anonymous):

So we chose 4 and 8 4(7b +8)(b -4)

OpenStudy (anonymous):

i ssee now